OpenStudy (anonymous):

What is 4/5 written as a decimal? A. 0.45 B. 0.54 C. 0.8 D. 1.25

sammixboo (sammixboo):

Let's convert 4/5 to 8/10.. They equal the same thing. Now do you know what is 8/10 as a decimal?

OpenStudy (michele_laino):

please note that: \[\frac{ 4 }{ 5 }=\frac{ 8 }{ 10 }\] so?

OpenStudy (anonymous):

The answer is C. right!?

sammixboo (sammixboo):

Correct!

OpenStudy (anonymous):

The denominator always tells you how many parts of 100 you are working with. So for example here, 5 means that we divide 100/5 = 20. Now the numerator tells you how many pieces of 20 you want... in this case 4. So, 4/5 = 80/100 = .8

OpenStudy (anonymous):

Yay!Thank You

OpenStudy (michele_laino):

@cjackson2002 that's right!

OpenStudy (anonymous):

0.8

OpenStudy (anonymous):

can i help you

OpenStudy (michele_laino):

@cjackson2002 Thank you!

OpenStudy (anonymous):

I have 5 more please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

7/25

OpenStudy (anonymous):

anymore

OpenStudy (anonymous):

what is the next problem

OpenStudy (anonymous):

http://www.asafraction.com/number/0.28